# Free Independence is not Definable

### v1

@@ 1. Introduction
The notion of a definable set is one of the central objects of study in classical model theory, often used in applications to other areas of mathematics (see for example [Pil00]). Generalizing the notion of a definable set to the setting of continuous model theory requires careful adjustment for the basic theory to transfer to the continuous setting, and the analysis of definable sets even in basic metric structures is far from complete. Previous works have focused on definable sets and functions in the Urysohn sphere and operators on Hilbert spaces, as well as the crucial notion of spectral gap in tracial von Neumann algebras [Gol12, Gol11, Gol23]. Additionally, a large number of basic objects of study in the theory of C-algebras are investigated through the lens of definability in [FHL+21, Chapter 3], and most recently minimal and maximal tensor norms were shown to be non-definable [GS25]. We continue this broad line of work by investigating the definability of the critical notion of free independence in both C-probability spaces and tracial von Neumann algebras.
Free independence is a central definition in the basic theory of free probability, introduced by Voiculescu in the 1980s. This theory has subsequently found numerous applications for understanding the structure of both von Neumann algebras and C-algebras.
Indeed, the best-known results on the question of isomorphism of the free group factors relies heavily on free probability [Răd92, Dyk94]. It has been also used to demonstrate structural properties of the corresponding C-algebras [HT05]. Recently, freeness has continued to play a role in C-algebra theory, contributing to a novel approach for demonstrating that certain algebras have nice regularity properties [AGEP24].
In this paper, we demonstrate that freeness is not definable, in the sense that the set of all pairs of elements $(x,y)$ which generate free subalgebras within a C-probability space (or a tracial von Neumann algebra) is not definable. The result follows from the characterisation of definable sets in terms of ultraproducts, and some important tools from operator algebras involving free complementation in ultrapowers. Crucially, we use Popa’s asymptotic freeness [Pop95] and Connes’ characterization of amenable factors [Con76].
1.1. Acknowledgements
@@ 2. Definable Sets in Continuous Logic
For completeness, we set out the basics of definable sets in continuous logic, and refer the reader to the exposition in [Gol23, Section 2] and [Har23, Section 8] for further information. For the benefit of the reader not versed in model theory, we remark that understanding the proofs that follow only relies upon the characterization given in equation (1), where $X$ is defined as in equation (2).
Definable sets are broadly those subsets that can be quantified over within a logical framework. For example, if a set $S$ is defined by a formula $\varphi$, then it is easy to express the statement “for all elements $x$ in SS,…” simply by writing “for all $x$, $\varphi(x)$ implies…” [^1] We are describing the case of classical (discrete) model theory here to give the basic flavour of definability. In the continuous setting, not every set defined by a formula is definable, see [FHL+21, Proposition 3.2.8].. Examples of such sets in unital C-algebras include the sets of self-adjoint elements, unitaries, and projections [FHL+21, Example 3.2.7]. Definable sets broaden this notion beyond just those sets exactly defined by formulas.
@@
  • $X(M)\subseteq Z(\phi^{M})$, where $Z(\phi^{M})$ denotes the zero-set in $M$ of the formula $\phi$, and
@@
In particular, it is not difficult to see that the second item above ensures that definable sets allow lifts from ultraproducts (see also [Har23, Theorem 8.2]):
@@ 3.1. Proof for $\mathrm{C}^{*}$-probability spaces
Since the universal UHF algebra $\mathcal{Q}$ is selfless ([Rob25, Theorem 5.2]), we may choose $x_{\infty}\in(\mathcal{Q})_{1}$ self-adjoint with spectrum $\sigma(x_{\infty})$ consisting of three points and a self-adjoint element $y\in(\mathcal{Q}_{\mathcal{U}})_{1}\setminus\mathbb{C}$ free from the constant sequence $x=[x_{\infty},\dots]\in\mathcal{Q}_{\mathcal{U}}$. Now suppose for a contradiction that freeness was definable, so that $X(\mathcal{Q}_{\mathcal{U}})=X(\mathcal{Q})^{\mathcal{U}}$, and so take two representative sequences $(x_{n})\subset\mathcal{Q}$ and $(y_{n})\subset\mathcal{Q}$ such that $x_{n}$ and $y_{n}$ are free, and $x_{n}]_{\mathcal{U}}=x$ and $y_{n}]_{\mathcal{U}}=y$ inside $\mathcal{Q}_{\mathcal{U}}$. We may further assume that $x_{n},y_{n}$ are self-adjoint for all $n\in\mathbb{N}$ by replacing $x_{n}$ by $(x_{n}+x_{n}^{*})/2$ and doing the same for $y_{n}$.
@@
Passing to the GNS-closure of $\mathcal{Q}$ with respect to its unique trace $\tau$ (which is isomorphic to the hyperfinite II1 factor $\mathcal{R}$), we have constructed two von Neumann subalgebras $A=\mathrm{W}^{*}(x_{n})$ and $B=\mathrm{W}^{*}(y_{n})$ which are free with respect to the unique trace $\tau$ on $\mathcal{R}$. Moreover, we have arranged that $\mathbb{C}^{3}$ embeds into $A$, and $B$ is not the complex numbers. In particular, we can find a copy of $\mathbb{C}^{2}\subseteq B$.
@@
Using Dykema’s [Dyk93, Theorem 2.3], one can express the free product as a direct sum of an interpolated free group factor $L(\mathbb{F}_{s})$ (see [Dyk94]) and a finite dimensional commutative von Neumann algebra. A calculation then shows that faithfulness of the traces implies that $s>1$ and so $(A,\tau|_{A})*(B,\tau|_{B})$ is non-amenable since it contains the non-amenable factor $L(\mathbb{F}_{s})$. ∎
@@ 3.2. Proof for tracial von Neumann algebras
The proof in the von Neumann setting is nearly identical - we use Popa’s asymptotic freeness result for $\mathcal{R}$ [Pop95] instead of selflessness of $\mathcal{Q}$ and have to work a bit harder to reach the conclusion of Lemma 3.2. We yet again choose $x_{\infty}\in(\mathcal{R})_{1}$ self-adjoint with $\sigma(x_{\infty})$ consisting of three points and choose $y\in(\mathcal{R}^{\mathcal{U}})_{1}\setminus\mathbb{C}$ self-adjoint and free from the constant sequence $x_{\infty},\dots]$ using [Pop95]. Assuming for a contradiction that freeness is definable, so that $X(\mathcal{R}^{\mathcal{U}})=X(\mathcal{R})^{\mathcal{U}}$, choose representative sequences $(x_{n})\subset\mathcal{R}$ and $(y_{n})\subset\mathcal{R}$ such that $x_{n}$ and $y_{n}$ are free, and $x_{n}]_{\mathcal{U}}=x$ and $y_{n}]_{\mathcal{U}}=y$ inside $\mathcal{R}^{\mathcal{U}}$. As before, we may further assume that $x_{n},y_{n}$ are self-adjoint for all $n\in\mathbb{N}$.
@@
We use [Kit85, Corollary 2] which states that for two normal Hilbert-Schmidt operators $a,b$ and any Lipschitz function $f\in C(\sigma(a)\cup\sigma(b))$ with Lipschitz constant $k$, we have the inequality
@@
We wish to apply the result to the sequence $x_{n}$ which converges to $x_{\infty}$ in the 2-norm. Since $x_{n}$ is bounded we can choose large enough $r\in\mathbb{R}_{+}$ such that $\sigma(x_{n})\cup\sigma(x_{\infty})$ is contained in the disk of radius $r$. Say $\sigma(x_{\infty})=\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ and fix disjointly supported Lipschitz functions $f_{1},f_{2},f_{3}$ such that $f_{i}$ is supported on a small neighbourhood of $\lambda_{i}$. As a consequence of 2-norm convergence, we then have
(5) $\int_{\sigma(x_{n})}f_{i}\mathop{}\!\mathrm{d}\mu_{x_{n}}$ $=\tau(f_{i}(x_{n}))$
(6) $\rightarrow\tau(f_{i}(x_{\infty}))=\int_{\sigma(x_{\infty})}f_{i}\mathop{}\!\mathrm{d}\mu_{x_{\infty}}\neq 0\,$,
where the measures appearing are the measures on $L^{\infty}(\sigma(x_{n}))$ induced by $\tau$. In particular we see that for $\mathcal{U}$-almost all $n$, the support of $f_{i}$ intersects $\sigma(x_{n})$ non-trivially and so the element $x_{n}$ must have at least three distinct points in its spectrum.
@@
The rest of the proof is identical to Section 3.1.

### v2

@@ 1. Introduction
The notion of a definable set is one of the central objects of study in classical model theory, often used in applications to other areas of mathematics (see for example [13]). Generalizing the notion of a definable set to the setting of continuous model theory requires careful adjustment for the basic theory to transfer to the continuous setting, and the analysis of definable sets even in basic metric structures is far from complete. Previous works have focused on definable sets and functions in the Urysohn sphere and operators on Hilbert spaces, as well as the crucial notion of spectral gap in tracial von Neumann algebras [8, 7, 9]. Additionally, a large number of basic objects of study in the theory of C-algebras are investigated through the lens of definability in [5, Chapter 3], and most recently minimal and maximal tensor norms were shown to be non-definable [6]. We continue this broad line of work by investigating the definability of the critical notion of free independence in both C-probability spaces and tracial von Neumann algebras.
Free independence is a central definition in the basic theory of free probability, introduced by Voiculescu in the 1980s. This theory has subsequently found numerous applications for understanding the structure of both von Neumann algebras and C-algebras. Indeed, the best-known results on the question of isomorphism of the free group factors relies heavily on free probability [15, 4]. It has been also used to demonstrate structural properties of the corresponding C-algebras [10]. Recently, freeness has continued to play a role in C-algebra theory, contributing to a novel approach for demonstrating that certain algebras have nice regularity properties [16, 1].
In this paper, we demonstrate that freeness is not definable, in the sense that the set of all pairs of elements $(x,y)$ which generate free subalgebras within a C-probability space (or a tracial von Neumann algebra) is not definable. The result follows from the characterisation of definable sets in terms of ultraproducts, and some important tools from operator algebras involving free complementation in ultrapowers. Crucially, we use Popa’s asymptotic freeness [14] and Connes’ characterization of amenable factors [2].
@@ 2. Definable Sets in Continuous Logic
For completeness, we set out the basics of definable sets in continuous logic, and refer the reader to the exposition in [9, Section 2] and [11, Section 8] for further information. For the benefit of the reader not versed in model theory, we remark that understanding the proofs that follow only relies upon the characterization of definable sets given by ultraproducts (see equation (1), where the definable set $X$ is defined as pairs of elements which are freely independent).
Definable sets are broadly those subsets that can be quantified over within a logical framework. For example, if a set $S$ is defined by a formula $\varphi$, then it is easy to express the statement “for all elements $x$ in SS,…” simply by writing “for all $x$, $\varphi(x)$ implies…” [^1] We are describing the case of classical (discrete) model theory here to give the basic flavour of definability. In the continuous setting, not every set defined by a formula is definable, see [5, Proposition 3.2.8].. Examples of such sets in unital C-algebras include the sets of self-adjoint elements, unitaries, and projections [5, Example 3.2.7]. Definable sets broaden this notion beyond just those sets exactly defined by formulas.
@@
  • $X(M)\subseteq Z(\phi^{M})$, where $Z(\phi^{M})$ denotes the zero-set in $M$ of the formula [^3] Here we mean usual formulas in continuous metric logic, although we could allow for definable predicates as well; see [5, Section 3.1]. $\phi$, and
@@
In particular, it is not difficult to see that the second item above ensures that definable sets allow lifts from ultraproducts (see also [11, Theorem 8.2]):
@@ 3. Non-definability of Freeness
We provide both proofs here in an elementary fashion; however, it should be noted that the C-algebraic result follows from the von Neumann algebraic result because the unique trace of a UHF algebra is a definable predicate in its theory, see [5, Theorem 3.5.5].
@@ 3.1. Proof for $\mathrm{C}^{*}$-probability spaces
Since the universal UHF algebra $\mathcal{Q}$ is selfless ([16, Theorem 5.2]), we may choose $x_{\infty}\in(\mathcal{Q})_{1}$ self-adjoint with spectrum $\sigma(x_{\infty})$ consisting of three points and a self-adjoint element $y\in(\mathcal{Q}_{\mathcal{U}})_{1}\setminus\mathbb{C}$ free from the constant sequence $x=[x_{\infty},\dots]\in\mathcal{Q}_{\mathcal{U}}$. Now suppose for a contradiction that freeness was definable, so that $X(\mathcal{Q}_{\mathcal{U}})=X(\mathcal{Q})^{\mathcal{U}}$, and so take two representative sequences $(x_{n})\subset\mathcal{Q}$ and $(y_{n})\subset\mathcal{Q}$ such that $x_{n}$ and $y_{n}$ are free, and $x_{n}]_{\mathcal{U}}=x$ and $y_{n}]_{\mathcal{U}}=y$ inside $\mathcal{Q}_{\mathcal{U}}$. We may further assume that $x_{n},y_{n}$ are self-adjoint for all $n\in\mathbb{N}$ by replacing $x_{n}$ by $(x_{n}+x_{n}^{*})/2$ and doing the same for $y_{n}$.
@@
Passing to the GNS-closure of $\mathcal{Q}$ with respect to its unique trace $\tau$ (which is isomorphic to the hyperfinite II1 factor $\mathcal{R}$), we have constructed two von Neumann subalgebras $A=\mathrm{W}^{*}(x_{n})$ and $B=\mathrm{W}^{*}(y_{n})$ which are free with respect to the unique trace $\tau$ on $\mathcal{R}$. Moreover, we have arranged that $\mathbb{C}^{3}$ embeds into $A$, and $B$ is not the complex numbers. In particular, we can find an isomorphic copy of $\mathbb{C}^{2}$ in $B$.
@@
Using Dykema’s [3, Theorem 2.3], one can express the free product as a direct sum of an interpolated free group factor $L(\mathbb{F}_{s})$ (see [4]) and a finite dimensional commutative von Neumann algebra. A calculation then shows that faithfulness of the traces implies that $s>1$ and so $(A,\tau|_{A})*(B,\tau|B)$ is non-amenable since it contains the non-amenable factor $L(\mathbb{F}_{s})$. ∎
@@ 3.2. Proof for tracial von Neumann algebras
The proof in the von Neumann setting is nearly identical - we use Popa’s asymptotic freeness result for $\mathcal{R}$ [14] instead of selflessness of $\mathcal{Q}$ and have to work a bit harder to reach the conclusion of Lemma 3.2. We yet again choose $x_{\infty}\in(\mathcal{R})_{1}$ self-adjoint with $\sigma(x_{\infty})$ consisting of three points and choose $y\in(\mathcal{R}^{\mathcal{U}})_{1}\setminus\mathbb{C}$ self-adjoint and free from the constant sequence $x_{\infty},\dots]$ using [14]. Assuming for a contradiction that freeness is definable, so that $X(\mathcal{R}^{\mathcal{U}})=X(\mathcal{R})^{\mathcal{U}}$, choose representative sequences $(x_{n})\subset\mathcal{R}$ and $(y_{n})\subset\mathcal{R}$ such that $x_{n}$ and $y_{n}$ are free, and $x_{n}]_{\mathcal{U}}=x$ and $y_{n}]_{\mathcal{U}}=y$ inside $\mathcal{R}^{\mathcal{U}}$. As before, we may further assume that $x_{n},y_{n}$ are self-adjoint for all $n\in\mathbb{N}$.
@@
We use [12, Corollary 2] which states that for two normal Hilbert-Schmidt operators $a,b$ and any Lipschitz function $f\in C(\sigma(a)\cup\sigma(b))$ with Lipschitz constant $k$, we have the inequality
@@
We wish to apply the result to the sequence $x_{n}$ which converges to $x_{\infty}$ in the 2-norm. Since $x_{n}$ is bounded we can choose large enough $r\in\mathbb{R}_{+}$ such that $\sigma(x_{n})\cup\sigma(x_{\infty})$ is contained in the disk of radius $r$. Say $\sigma(x_{\infty})=\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ and fix disjointly supported Lipschitz functions $f_{1},f_{2},f_{3}\in C(\mathbb{R},\mathbb{R}_{+})$ such that $f_{i}$ is only supported on a small neighbourhood of $\lambda_{i}$. As a consequence of 2-norm convergence, we then have for each $i$,
[…]
where the measures $\mu_{x_{n}}$ are the measures on $L^{\infty}(\sigma(x_{n}))$ induced by $\tau$. In particular we see that for $\mathcal{U}$-almost all $n$, the support of $f_{i}$ intersects $\sigma(x_{n})$ non-trivially and so the element $x_{n}$ must have at least three distinct points in its spectrum.
@@
The rest of the proof is analogous to the proof given in Section 3.1.